OpenStudy (anonymous):

Find the unique polynomial in factored form with real coefficients that meets these cond: degree:4 zeros:-5,12-3i,f(0)=130

OpenStudy (anonymous):

fx= (x+5)^2(x-1)(x-(2-3i))+130 I got this

OpenStudy (anonymous):

alternate forms assuming f and x are real. fx = x4 + 7 x3 - 3 x2 + i ( 3 x3 + 27 x2 + 45 x - 75) - 55 x + 180

OpenStudy (dumbcow):

not quite ... with complex zeros come in pairs 2+3i is also a zero and dont add 130 at end...only if x=0, f = 130 there is an "a" coefficient in fron of factors dictated by point f(0) = 130 \[f(x) = a(x+5)(x-1)(x-(2-3i))(x-(2+3i))\] you have to expand to get rid of "i" and solve for "a"

OpenStudy (anonymous):

wait but where are the exponents and 130.. ? you don't add that in? @dumbcow

OpenStudy (dumbcow):

no for example ... say zeros are -5,1 f(0) = 130 you would say... f(x) = (x+5)(x-1) +130 but if you plug in x=0, it wont work f(0) = (5)(-1) +130 = 125 make sense?

OpenStudy (anonymous):

oh I see!

OpenStudy (anonymous):

So how can I approach this problem

OpenStudy (dumbcow):

write it in factored form and put a coefficient in front ...in prev example --> f(x) = a(x+5)(x-1) f(0) = a(5)(-1) = 130 -5a = 130 a = -26

OpenStudy (anonymous):

ohhhh wait can we totally disregard the imaginary numbers

OpenStudy (dumbcow):

now for complex zeros ...you can write it out as factors then expand but i suggest this approach \[x = 2-3i\] work backwards until right side is 0 \[x-2 = -3i\] \[(x-2)^{2} = (-3i)^{2}\] \[x^{2} -4x +4 = -9\] \[x^{2} -4x +13 = 0\] that is the factor for both complex zeros \[f(x) = a(x+5)(x-1)(x^{2} -4x+13)\]

OpenStudy (anonymous):

iohhh

OpenStudy (anonymous):

a=-26

OpenStudy (dumbcow):

haha no i only did that to make it easier to understand the "a" concept no a is not -26 ... just for my example

OpenStudy (anonymous):

lol oh ok

OpenStudy (dumbcow):

its not hard solving for a though ... just plug in x=0

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

wait and y=130?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

oh wow so

OpenStudy (anonymous):

y=2(x+5)(x-1)(x^2-4x+13)

OpenStudy (dumbcow):

yep...well it should be -2 in front

OpenStudy (anonymous):

YESYES lol sorry that's what I got too ahaha

OpenStudy (anonymous):

THNX!

OpenStudy (dumbcow):

yw :)